# Some Bayesian Perspectives on Clinical Trials

### v1

@@ Abstract
We provide a Bayesian perspective on three interconnected aspects of clinical trial design: prior specification, sequential adaptive allocation, and decision-theoretic optimization. For prior specification, we argue that treatment effects in clinical trials are known a priori to be small, rendering default noninformative priors such as Jeffreys’ prior inappropriate; priors calibrated to historical effect sizes or LD50 relationships are both more honest and more efficient. For sequential design, we show how Thompson’s (1933) probability-matching rule connects to modern adaptive randomization, and how backward induction on sufficient statistics—following Christen and Nakamura [2003] and Carlin et al. [1998]—reduces the seemingly intractable infinite-horizon stopping problem to a finite table. For trial optimization, we review the utility-based framework of Thall and Cook [2004] that jointly models efficacy and toxicity, enabling dose-finding designs that maximize patient benefit rather than merely controlling error rates. We illustrate these ideas through the ECMO trial, the CALGB 49907 breast cancer trial, and modern platform trials, and discuss the 2026 FDA draft guidance on Bayesian methodology.
@@ 1 Introduction
The standard paradigm for clinical trial design—fix the sample size, randomize equally, analyze once at the end—was built to control the probability of approving an ineffective drug. It succeeded at that goal. But the framework treats patients enrolled during the trial as expendable inputs to a statistical procedure whose conclusions matter only afterward. If early data show one treatment to be clearly inferior, patients continue receiving it. If the trial is underpowered, patients’ participation generates inconclusive evidence.
@@
Bayesian methods offer a different organizing principle: probability is a measure of uncertainty, updated coherently as data arrive. This principle has three consequences for clinical trials that we develop in this paper:
  1. 1.
@@
    Priors encode what is already known. Treatment effects in clinical trials are small—typically a few percentage points of absolute risk reduction. This prior knowledge, ignored by noninformative analyses, can improve both efficiency and calibration when formalized correctly.
  2. 2.
@@
    Sequential decisions can be made optimally. When utility depends on sufficient statistics, the infinite-horizon stopping problem reduces to a computable table via backward induction. Combined with Thompson’s (1933) probability-matching rule, this yields adaptive designs that balance learning against treating.
  3. 3.
@@
We are not the first to advocate Bayesian clinical trials. Novick and Grizzle [1965] introduced Bayesian analysis of clinical data using conjugate priors for categorical outcomes. Berry et al. [2010] provide a comprehensive treatment of Bayesian adaptive methods. Berry [2025] reviews the state of the art. Our contribution is to connect three specific arguments—that default priors are wrong, that backward induction is tractable, and that utility-based design unifies ad hoc adaptations—and to illustrate them through trials where these ideas made a material difference.
Section 2 discusses prior specification. Section 3 covers sequential adaptive designs. Section 4 develops the decision-theoretic framework. Section 5 presents case studies. Section 6 discusses the regulatory landscape. Section 7 concludes.
@@ 2.1 Why default priors fail for treatment effects
Jeffreys’ prior for the binomial proportion is $\text{Beta}(1/2,1/2)$, which places substantial mass near 0 and 1. Under independent Jeffreys priors for $\theta_{1}$ and $\theta_{0}$, the implied prior on $\delta$ is U-shaped, concentrating mass near $\pm 1$. This contradicts everything known about drug development: treatment effects are small. Among drugs that reach Phase III, the median absolute risk reduction is on the order of 5–10 percentage points [Berry, 2012]. A prior that places most mass on treatment effects near $\pm 100\%$ is not noninformative—it is misinformative.
The uniform prior $\text{Beta}(1,1)$ is less extreme but still problematic. It implies that a response rate of 95% is as likely as 50% a priori, and that the treatment effect $\delta$ has prior standard deviation of approximately 0.41—far larger than any plausible effect in most therapeutic areas.
@@
Novick and Grizzle [1965] recognized this issue and proposed “logical probability” priors calibrated to prior data. Their natural conjugate Bayes densities express prior information as equivalent to $(m,z)$ prior observations with $z$ successes, so the prior $\text{Beta}(z+1,m-z+1)$ has the interpretation of $m$ prior patients. This parameterization makes the prior’s influence transparent: it adds $m$ pseudo-observations to the data. The effective sample size concept was later formalized by Morita et al. [2008].
@@ 2.2 Informative priors from historical data
A more principled approach uses historical data to construct informative priors. Racine et al. [1986] demonstrated this systematically across four areas of pharmaceutical research at Ciba-Geigy: LD50 experiments, crossover trials, bioequivalence assessment, and population pharmacokinetics. For LD50 experiments, where small sample sizes cause fiducial intervals to degenerate to the entire real line, Bayesian priors calibrated to toxicity class boundaries produced posterior densities that were both well-defined and practically useful. For population pharmacokinetic models—hierarchical nonlinear random effects models where individual parameters $\theta_{i}$ vary across patients—the Bayesian framework borrows strength across individuals, enabling prediction of a new patient’s drug concentration profile from as few as two blood draws when combined with population data. If a drug’s mechanism of action is understood, prior knowledge about the control arm response rate is often substantial. For dose-response relationships, the LD50 (the dose producing 50% response) provides a natural anchor: a prior on the dose-response curve centered at the LD50 with variance reflecting uncertainty about the curve’s slope encodes genuine pharmacological knowledge.
A fundamental difficulty in specifying priors for dose-response models is that regression coefficients are abstract and model-dependent: the meaning of $\beta$ in a logistic model differs from its meaning in a probit model, so eliciting priors directly on $\beta$ is unnatural. Bedrick et al. [1996] resolved this by introducing conditional means priors: the investigator specifies prior distributions on observable quantities—the probability of response at selected covariate values—and the prior on $\beta$ is induced through the link function. For a binomial regression with $p$ predictors, one specifies independent $\text{Beta}(a_{1i},a_{2i})$ priors on the success probabilities $\tilde{m}_{i}$ at $p$ chosen covariate vectors $\tilde{\mathbf{x}}_{i}$, and the induced prior on $\beta$ takes the form
@@
where $F$ is the CDF associated with the link function and $f$ its density. The approach is concrete. In the Challenger O-ring analysis, Bedrick et al. [1996] modeled the probability that any O-ring fails as a function of temperature via $g(m_{i})=\beta_{0}+\beta_{1}(\tau_{i}-\bar{\tau})$ and specified independent beta priors on the failure probabilities at two temperatures: $\tilde{m}_{1}\sim\text{Beta}(1,0.577)$ at $55^{\circ}$F (giving $\Pr[\tilde{m}_{1}>1/2]=2/3$, reflecting a belief that failure is likely at low temperatures) and $\tilde{m}_{2}\sim\text{Beta}(0.577,1)$ at $75^{\circ}$F (giving $\Pr[\tilde{m}_{2}<1/2]=2/3$). These priors on observable failure rates are easy to elicit; the induced prior on $(\beta_{0},\beta_{1})$ via (1) then depends on the link function, so the same scientific beliefs produce different coefficient priors for logistic versus probit models without requiring the investigator to think about model-dependent abstractions. The idea traces back to Racine et al. [1986], who used independent beta priors on two success probabilities to induce bivariate priors on probit regression coefficients in bioequivalence studies at Ciba-Geigy.
@@
For logistic regression specifically, the induced prior is equivalent to a data augmentation prior—pseudo-observations at the design points with fractional sample sizes—connecting directly to the Pólya-Gamma scheme below. Hanson et al. [2014] extended this through informative $g$-priors for logistic regression, where the investigator specifies only a prior on the overall population success probability and the $g$-prior structure induces a proper joint prior on all coefficients. Together with Chaloner [1993], who characterized optimal Bayesian designs for dose-response experiments under such priors, these methods give the trialist a complete toolkit: specify beliefs about response rates at clinically meaningful doses, let the link function induce priors on model parameters, and design the experiment to maximize information about the LD50 or other quantities of interest.
The FDA’s 2026 draft guidance on Bayesian methodology [Food and Drug Administration, 2026] identifies several settings where informative priors from historical data have supported regulatory submissions: borrowing from previous clinical trials of the same drug, augmenting concurrent controls with external data, pediatric extrapolation from adult trials, and borrowing across similar disease subtypes. In each case, the prior’s influence must be quantified—for example, through its effective sample size—and sensitivity analyses must demonstrate robustness to prior-data conflict.
Dynamic borrowing methods, where the degree of information transfer depends on observed consistency between historical and current data, mitigate the risk of static priors that do not respond to conflict. The Bayesian hierarchical framework provides a natural mechanism: the between-study variance parameter $\tau^{2}$ governs borrowing, with large $\tau^{2}$ producing skeptical priors that are quickly overwhelmed by current data.
2.3 Computational advances: Pólya-Gamma augmentation

### v2

@@ Abstract
We examine three landmark clinical trials—ECMO, CALGB 49907, and I-SPY 2—through a unified Bayesian framework connecting prior specification, sequential adaptation, and decision-theoretic optimisation. For ECMO, the posterior probability of treatment superiority is robust across the range of priors examined. For CALGB, predictive probability monitoring stopped enrolment at 633 instead of 1800 patients. For I-SPY 2, adaptive enrichment graduated nine of 23 arms to Phase III. These case studies motivate a methodological contribution: exact backward induction for two-arm binary trials, where Beta-Binomial conjugacy yields closed-form transitions on the integer lattice of success counts with no quadrature. A Pólya-Gamma augmentation bridges this to covariate-adjusted logistic regression. Simulation reveals a fundamental tension: the optimal Bayesian design reduces expected sample sizes to 14–26 per arm (versus 42–100 for alternatives) but with substantially lower power. A calibrated variant embedding the declaration threshold in the terminal utility improves power while maintaining sample-size savings; varying the per-stage cost traces a power frontier for selecting the preferred operating point, with suitability highest in patient-sparing contexts such as rare diseases and paediatrics. The Pólya-Gamma Laplace approximation is validated against exact calculations (mean absolute error below 0.01). We discuss implications for the 2026 FDA draft guidance on Bayesian methodology.
Keywords: adaptive randomisation; backward induction; Bayesian clinical trials; Beta-Binomial; Pólya-Gamma augmentation; predictive probability; prior elicitation; Thompson sampling.
@@ 1 Introduction
The standard paradigm for clinical trial design—fix the sample size, randomise equally, analyse once at the end—was built to control the probability of approving an ineffective drug. But the framework prioritises error-rate control at the expense of within-trial patient outcomes: patients enrolled during the trial contribute to a statistical procedure whose conclusions are drawn only afterward. If early data show one treatment to be clearly inferior, patients continue receiving it. If the trial is underpowered, patients’ participation generates inconclusive evidence.
The inefficiency is quantifiable. Consider a trial testing a new antihypertensive drug against standard care, with blood pressure reduction as the primary endpoint. The classical power calculation requires fixing the sample size to detect the minimum clinically important difference—say, 5 mmHg—with 90% power at the two-sided 5% level. With standard deviation $\sigma=15$ mmHg, this yields approximately 190 patients per group, or 380 total. But suppose Phase II data suggest the true effect is likely around 10 mmHg, not 5 mmHg, with uncertainty reflected in a prior distribution $\delta\sim N(10,5^{2})$. A Bayesian design that formally incorporates this prior information requires fewer patients to achieve comparable expected power (averaged over the prior on $\delta$), because the prior concentrates mass on plausible effect sizes rather than requiring protection against the worst case (Inoue et al., 2005). The classical calculation ignores what is already known from Phase II; the Bayesian calculation uses it.
The divergence reflects a deeper conceptual difference. The frequentist framework asks: “If the null hypothesis were true and this experiment were repeated many times, would the decision rule produce false positives in fewer than 5% of repetitions?” The question is about error rates across hypothetical trials that will never be run. The Bayesian framework asks: “Given the data actually observed and the prior information available, what do we believe about the treatment effect, and what decision maximises expected utility?” When trials are monitored sequentially, this distinction becomes operational. The frequentist solution—alpha-spending functions that allocate type-I error across interim analyses—imposes a statistical penalty for looking at the data: to maintain overall $\alpha=0.05$, early stopping boundaries require inflated critical values at initial looks (for example, $z\approx 4$ under O’Brien–Fleming spending), which can delay stopping even when evidence favours one arm. Futility boundaries partly mitigate this, but the constraint remains that each additional look consumes part of the error budget. The Bayesian posterior, by contrast, updates after every observation without a multiplicity penalty in the likelihood, because the stopping rule enters through the sampling model, not through the inference. This does not eliminate the need for calibration: Bayesian designs intended for confirmatory use still require pre-specified decision thresholds and simulation-based verification of operating characteristics (Section 3.7). Stopping rules based on posterior or predictive probabilities answer the question trialists actually face: should we stop now, or is the expected value of continuing positive?
The philosophical divide does not preclude practical agreement at the design stage—and the agreement is deeper than most practitioners realise. The frequentist formula for sample size determination is among the most familiar calculations in biostatistics. For a one-sided test of $H_{0}\colon\theta=\theta_{0}$ versus $H_{1}\colon\theta=\theta_{1}$ ($\theta_{1}>\theta_{0}$) with Normal data and known variance $\sigma^{2}$:
$n_{F}=(z_{\alpha}+z_{\beta})^{2}\left(\frac{\sigma}{\delta}\right)^{2},$ (1)
where $\delta=\theta_{1}-\theta_{0}$ is the target difference, $z_{\alpha}=\Phi^{-1}(1-\alpha)$, and $1-\beta$ is the desired power. Every trialist knows this formula. What is less widely appreciated is that a Bayesian who frames the same question differently—seeking the smallest $n$ such that the expected utility of the experiment exceeds a threshold—arrives at the same sample size, not just for one $\delta$ but for all $\delta$ simultaneously.
Inoue et al. (2005) formalised this correspondence by recasting sample-size determination as a decision problem with an explicit utility—a goal function—rather than conditional error rates. The Bayesian does not ask “what are the type I and type II error rates?” but “what sample size minimises the expected cost of wrong decisions?” Specifically, the Bayesian assigns prior probability $\pi$ to $H_{0}$ and $1-\pi$ to $H_{1}$, and adopts a 0-11-$K$ loss: no cost for a correct decision, cost $1$ for failing to reject $H_{0}$ when $H_{1}$ is true, and cost $K$ for rejecting $H_{0}$ when it is true. Minimising posterior expected loss yields a decision rule: do not reject $H_{0}$ whenever the posterior probability of $H_{0}$ exceeds $1/(1+K)$. For Normal data with known $\sigma^{2}$, this reduces to the cut-off
$\bar{x}\;\leq\;\frac{\sigma^{2}\log\bigl(K\,\pi/(1-\pi)\bigr)}{n\,\delta}\;+\;\frac{\theta_{1}+\theta_{0}}{2}\,.$
The goal function—equivalently, the expected utility of the experiment—is the utility-weighted probability of a correct decision (reducing to a pure classification rate when $K=1$):
$G_{B}(n)$ $=K\pi\,P_{\theta_{0}}(\text{accept}\;H_{0})+(1-\pi)\,P_{\theta_{1}}(\text{reject}\;H_{0})$
$=K\pi\,\Phi\!\left(\frac{\sigma\log\bigl(K\,\pi/(1\!-\!\pi)\bigr)}{\sqrt{n}\,\delta}+\frac{\delta\sqrt{n}}{2\sigma}\right)+(1\!-\!\pi)\,\Phi\!\left(\frac{\delta\sqrt{n}}{2\sigma}-\frac{\sigma\log\bigl(K\,\pi/(1\!-\!\pi)\bigr)}{\sqrt{n}\,\delta}\right),$ (2)
where the second line follows by standardising $\bar{x}$ under each hypothesis. The Bayesian finds the smallest $n$ such that $G_{B}(n)\geq r^{*}$, a target level of expected utility. This is the Bayesian analogue of the frequentist power calculation—but framed entirely in terms of expected utility rather than conditional error rates. In this canonical setting, favourable type I and type II error rates emerge as a byproduct of maximising expected utility rather than as the primary design target.
Substituting $n=n_{F}$ from (1) into (1), every occurrence of $\delta$ and $\sigma$ cancels:
$G_{B}(n_{F})=K\pi\,\Phi\!\left(\frac{\log(K\pi/(1\!-\!\pi))}{|z_{\alpha}+z_{\beta}|}+\frac{|z_{\alpha}+z_{\beta}|}{2}\right)+(1\!-\!\pi)\,\Phi\!\left(\frac{|z_{\alpha}+z_{\beta}|}{2}-\frac{\log(K\pi/(1\!-\!\pi))}{|z_{\alpha}+z_{\beta}|}\right).$ (3)
The right-hand side depends only on $(\alpha,\beta,\pi,K)$, not on $\delta$ or $\sigma$. With $\pi=1/2$ and $K=1$ (symmetric prior and loss), the expression simplifies to $\Phi\bigl((z_{\alpha}+z_{\beta})/2\bigr)$, which at $\alpha=0.05$ and $\beta=0.10$ gives $r^{*}=0.928$. A frequentist who calculates $n_{F}=857$ for $\sigma=1$, $\delta=0.10$ and a Bayesian who requires $r^{*}=0.928$ will agree on n=857n=857—and they will continue to agree at $\delta=0.05$ (n=3,426n=3{,}426), $\delta=0.20$ ($n=214$), and at every other $\delta$. The choice of $\delta$ in the frequentist calculation implicitly encodes the same information as the Bayesian prior: both determine the signal-to-noise ratio at which the study becomes conclusive.
The result extends to composite hypotheses $H_{0}\colon\theta\leq\theta_{0}$ versus $H_{1}\colon\theta>\theta_{0}$ under a Normal prior $\theta\sim N(\theta_{0},\tau^{2})$, provided $\tau^{2}=C\delta^{2}$ for a constant C>0C>0—that is, the prior standard deviation scales with the target effect. It also extends to Lindley information as the Bayesian goal, where the information target $\mathcal{I}^{*}=\tfrac{1}{2}\log(1+C(z_{\alpha}+z_{\beta})^{2})$ is again constant in $\delta$, and to Bernoulli data, where the correspondence holds approximately.
The practical implication is that, in these canonical sample-size settings, Bayesian and frequentist approaches are not in conflict—they are different parametrisations of the same calculation. Every trialist knows the power formula (1), but few recognise that the familiar inputs $(\alpha,\beta,\delta)$ implicitly encode a prior and a utility: the “relevant difference” $\delta$ plays the role of the prior location, and the error rates $\alpha$ and $\beta$ map to the asymmetric loss $K$ and classification threshold $r^{*}$. The power calculation is not wrong—it is a goal function in disguise.
The tension we develop in this paper is therefore not between Bayesian and frequentist philosophies per se, but between designs that optimise expected utility and designs that optimise frequentist operating characteristics. At the design stage, as (3) shows for the canonical Normal setting with known variance, the two coincide. It is in sequential monitoring that they diverge. The frequentist who monitors a trial sequentially cannot simply look at the accumulating data: each interim analysis consumes part of the type I error budget, the alpha-spending function must be pre-specified, and modifying it mid-trial requires a protocol amendment. In practice, this means that even when early data strongly favour one arm, the frequentist framework imposes a statistical cost for having looked. The Bayesian posterior, by contrast, updates after every observation without such a penalty. In the sequential setting, the organising principle remains expected utility—not type I and type II error—and the goal function (1) generalises naturally by conditioning on the data observed so far. This asymmetry—the freedom to inspect data at any time, optimising expected utility at each step—is the operational advantage that motivates the sequential designs we develop next.
@@
The foundational idea is simple: probability is a measure of uncertainty, updated coherently as data arrive. This principle has four consequences for clinical trials that we develop in this paper:
  1. (a)
@@
    Priors encode what is already known. Treatment effects in clinical trials are small—typically a few percentage points of absolute risk reduction. This prior knowledge, ignored by noninformative analyses, can improve both efficiency and calibration when formalised correctly. More broadly, the posterior precision at any sample size equals the sum of prior precision and data precision. With an informative prior, the posterior starts concentrated and tightens with each observation; without one, the first observations are wasted re-learning what was already known.
  2. (b)
    Small samples suffice for strong conclusions. Because the posterior accumulates evidence continuously, a trial can stop as soon as the posterior probability of superiority (or futility) crosses a decision threshold—whether that takes 12 patients or 600. The Bayesian framework imposes no likelihood-based penalty for early inspection—the posterior is valid regardless of the stopping rule—so evidence that arrives quickly is acted on quickly. (The frequentist operating characteristics of repeated monitoring still require calibration; see Section 3.7.) With an informative prior, fewer observations are needed to cross any given threshold, because the prior has already contributed information equivalent to its effective sample size.
  3. (c)
@@
    Sequential decisions can be made optimally. When utility depends on sufficient statistics, the stopping problem reduces to a computable table via backward induction. Combined with Thompson’s (1933) probability-matching rule, this yields adaptive designs that balance learning against treating.
  4. (d)
@@
The practical upshot is that Bayesian trials can reach definitive conclusions—in either direction—with fewer patients than fixed-sample designs, though the magnitude of the saving depends on the utility function and comes with trade-offs in power and type I error that we quantify in Section 3.7. Informative priors amplify the sample-size advantage. The case studies in this paper illustrate the point repeatedly: ECMO reached a strong conclusion with 12 patients (Section 5), CALGB 49907 stopped 65% early, and a hierarchical model applied to multi-centre data produces a 98.5% posterior probability of treatment efficacy from centres with as few as five patients per arm (Section 2).
We are not the first to advocate Bayesian clinical trials. Novick and Grizzle (1965) introduced Bayesian analysis of clinical data using conjugate priors for categorical outcomes. Berry et al. (2010) provide a comprehensive treatment of Bayesian adaptive methods. Berry (2025) reviews the state of the art. Our contributions are threefold. First, we develop exact backward induction for two-arm binary trials, exploiting the Beta-Binomial conjugacy to obtain closed-form transition probabilities on the integer lattice of success counts. Unlike the Normal model of Brockwell and Kadane (2003), which requires numerical quadrature on a continuous grid, the binary model yields exact optimal stopping rules with no discretisation error. A calibrated variant embeds the declaration threshold directly in the terminal utility, yielding a power frontier that makes the sample-size–power trade-off explicit and tunable. Second, we show how Pólya-Gamma augmentation (Polson et al., 2013) bridges this framework to logistic regression with covariate adjustment, and validate the resulting Laplace approximation against exact Beta posterior calculations. Third, we demonstrate through three landmark trials how the Bayesian principles of prior specification, sequential adaptation, and decision-theoretic optimisation interact in practice: informative priors sharpen inference when data are sparse (ECMO), predictive probabilities enable early stopping that saves hundreds of patients (CALGB 49907), and utility-based adaptive enrichment accelerates drug development in heterogeneous populations (I-SPY 2). For the ECMO trial, we provide a dedicated backward induction re-analysis with asymmetric historical priors.
Section 2 discusses prior specification. Section 3 covers sequential adaptive designs, including exact binary backward induction, the Pólya-Gamma bridge, and a simulation study with power frontier analysis. Section 4 develops the decision-theoretic framework. Section 5 presents case studies, including a backward induction re-analysis of the ECMO trial. Section 6 discusses the regulatory landscape. Section 7 concludes.
@@ 2.1 Why default priors fail for treatment effects
Jeffreys’ prior for the binomial proportion is $\text{Beta}(1/2,1/2)$, which places substantial mass near 0 and 1. Under independent Jeffreys priors for $\theta_{1}$ and $\theta_{0}$, the implied prior on $\delta$ is U-shaped, concentrating mass near $\pm 1$. This contradicts everything known about drug development: treatment effects are small. Among drugs that reach Phase III, absolute risk reductions are typically modest—often on the order of a few percentage points (Berry, 2012). A prior that places most mass on treatment effects near $\pm 100\%$ is not noninformative—it is misinformative.
The uniform prior $\text{Beta}(1,1)$ is less extreme but still problematic. It implies that a response rate of 95% is as likely as 50% a priori, and that the treatment effect $\delta$ has prior standard deviation of approximately 0.41—far larger than any plausible effect in most therapeutic areas. From a regularisation perspective, the choice of prior is a modelling decision with direct inferential consequences: a uniform prior corresponds to maximum likelihood, while informative priors provide the Bayesian analogue of penalised estimation (Polson and Sokolov, 2019).
@@
Novick and Grizzle (1965) recognised this issue and proposed “logical probability” priors calibrated to prior data. Their natural conjugate Bayes densities express prior information as equivalent to $(m,z)$ prior observations with $z$ successes, so the prior $\text{Beta}(z+1,m-z+1)$ has the interpretation of $m$ prior patients. This parametrisation makes the prior’s influence transparent: it adds $m$ pseudo-observations to the data. The effective sample size concept was later formalised by Morita et al. (2008).
@@ 2.2 Informative priors from historical data
A more principled approach uses historical data to construct informative priors. Racine et al. (1986) demonstrated this systematically across four areas of pharmaceutical research at Ciba-Geigy: LD50 (median lethal dose) experiments, crossover trials, bioequivalence assessment, and population pharmacokinetics. For LD50 experiments, where small sample sizes cause fiducial intervals to degenerate to the entire real line, Bayesian priors calibrated to toxicity class boundaries produced posterior densities that were both well-defined and practically useful. For population pharmacokinetic models—hierarchical nonlinear random effects models where individual parameters $\theta_{i}$ vary across patients—the Bayesian framework borrows strength across individuals, enabling prediction of a new patient’s drug concentration profile from as few as two blood draws when combined with population data. If a drug’s mechanism of action is understood, prior knowledge about the control arm response rate is often substantial. For dose-response relationships, the LD50 (the dose producing 50% response) provides a natural anchor: a prior on the dose-response curve centred at the LD50 with variance reflecting uncertainty about the curve’s slope encodes genuine pharmacological knowledge.
A fundamental difficulty in specifying priors for dose-response models is that regression coefficients are abstract and model-dependent: the meaning of $\beta$ in a logistic model differs from its meaning in a probit model, so eliciting priors directly on $\beta$ is unnatural. Bedrick et al. (1996) resolved this by introducing conditional means priors: the investigator specifies prior distributions on observable quantities—the probability of response at selected covariate values—and the prior on $\beta$ is induced through the link function. For a binomial regression with $p$ predictors, one specifies independent $\text{Beta}(a_{1i},a_{2i})$ priors on the success probabilities $\tilde{m}_{i}$ at $p$ chosen covariate vectors $\tilde{\mathbf{x}}_{i}$, and the induced prior on $\beta$ takes the form
@@
where $F$ is the CDF associated with the link function and $f$ its density. In the Challenger O-ring analysis, Bedrick et al. (1996) modelled the probability that any O-ring fails as a function of temperature via $g(m_{i})=\beta_{0}+\beta_{1}(\tau_{i}-\bar{\tau})$ and specified independent beta priors on the failure probabilities at two temperatures: $\tilde{m}_{1}\sim\text{Beta}(1,0.577)$ at $55^{\circ}$F (giving $\Pr[\tilde{m}_{1}>1/2]=2/3$, reflecting a belief that failure is likely at low temperatures) and $\tilde{m}_{2}\sim\text{Beta}(0.577,1)$ at $75^{\circ}$F (giving $\Pr[\tilde{m}_{2}<1/2]=2/3$). These priors on observable failure rates are easy to elicit; the induced prior on $(\beta_{0},\beta_{1})$ via (4) then depends on the link function, so the same scientific beliefs produce different coefficient priors for logistic versus probit models without requiring the investigator to think about model-dependent abstractions. Racine et al. (1986) used a similar idea informally, specifying independent beta priors on two success probabilities to induce bivariate priors on probit regression coefficients in bioequivalence studies at Ciba-Geigy; Bedrick et al. (1996) formalised and generalised the approach.
@@
For logistic regression specifically, the induced prior is equivalent to a data augmentation prior—pseudo-observations at the design points with fractional sample sizes—connecting directly to the Pólya-Gamma scheme below. Hanson et al. (2014) extended this through informative $g$-priors for logistic regression, where the investigator specifies only a prior on the overall population success probability and the $g$-prior structure induces a proper joint prior on all coefficients. Together with Chaloner (1993), who characterised optimal Bayesian designs for dose-response experiments under such priors, these methods enable the trialist to specify beliefs about response rates at clinically meaningful doses, let the link function induce priors on model parameters, and design the experiment to maximise information about the LD50 or other quantities of interest.
The FDA’s 2026 draft guidance on Bayesian methodology (Food and Drug Administration, 2026) identifies several settings where informative priors from historical data have supported regulatory submissions: borrowing from previous clinical trials of the same drug, augmenting concurrent controls with external data, paediatric extrapolation from adult trials, and borrowing across similar disease subtypes. In each case, the prior’s influence must be quantified—for example, through its effective sample size—and sensitivity analyses must demonstrate robustness to prior-data conflict.
Dynamic borrowing methods, where the degree of information transfer depends on observed consistency between historical and current data, mitigate the risk of static priors that do not respond to conflict. The Bayesian hierarchical framework provides a natural mechanism: the between-study variance parameter $\tau^{2}$ governs borrowing, with large $\tau^{2}$ producing sceptical priors that are quickly overwhelmed by current data.
2.3 Pólya-Gamma augmentation